OpenStudy (anonymous):

The sum of four times a number, and 8, is equal to three times the number? In a equation. 4x+8=3x?

OpenStudy (dtan5457):

Yeah, that looks right

OpenStudy (dtan5457):

Now just solve for x

OpenStudy (dtan5457):

Which is..?

OpenStudy (anonymous):

-8

OpenStudy (dtan5457):

Yep

OpenStudy (dtan5457):

And you can even check because -8 times 4 +8=-24 -8 times 3=-24 -8 is correct